OpenStudy (anonymous):

A medal for who can answer the following: Helena made trail mix and wants to list the ingredients in order of increasing weight. She used one and five eighths lb of mixed nuts, five eighths lb of chocolate covered raisins, and one and two eighths lb of dried mixed fruit. If she decides to include one and seven eighths lb of sunflower seeds, which list shows the order of ingredients from least weight to greatest weight? seeds, nuts, raisins, fruit nuts, seeds, raisins, fruit raisins, fruit, nuts, seeds raisins, fruit, seeds, nuts

OpenStudy (anonymous):

4 th one

OpenStudy (anonymous):

Raisins, Fruit, Nuts, seeds 5/8,1 2/8,1 5/8, 1 7/8=5/8,5/4,13/8,15/8=0.625,1.25,1.625,1.875